OpenStudy (anonymous):

Ellen, a resident of Metropolis, pays Metropolis an annual tax of $50 plus 1.6% of her annual income. If Ellen paid $1,298 in tax, what was Ellen’s income?

OpenStudy (anonymous):

50 + 1.6/100 of x =1298 can u find x from here

OpenStudy (mathstudent55):

First subtract $50 from the tax she paid. Then the remaining amount is 1.6% of her salary. Call it T for tax. If T is 1.6% of S, the salary, then 0.016T = S Plug in T and calculate S.

OpenStudy (anonymous):

um email me im confused spareb665@yahoo.com

OpenStudy (mathstudent55):

Can you start with $1298 and subtract $50 from that? What do you get?

OpenStudy (anonymous):

1248?

OpenStudy (mathstudent55):

Right. Now $1248 is 1.6% of her salary, right?

OpenStudy (anonymous):

he will get 0.016t=1248 for t?

OpenStudy (mathstudent55):

The question now is: $1248 is 1.6% of what number? Translated into math this is: 1248 = 1.6% * x In math, "of" usually means multiplication; is means =, percen means 1/100. 1248 = (1.6/100)x or 0.016x = 1248 Divide both sides by 0.016 to get x, her salary.

OpenStudy (anonymous):

78000? coprect?

OpenStudy (mathstudent55):

Correct. Good job!